=== 2019/3/4(Mon.) ===

=== Schedule ===
 * Center
  * --(2/26-27: Venting)--
  * 2/28-3/5: Fixing PRMs and SRMs
  * 3/2-5: Re-alignment
  * 3/6-7: Closing

 * Yend
  * --(1/25-2/20: Yend Pcal finalization)-- 
  * 2/25-3/4: Yend WAB finalization
  * 3/5-6: Closing
  * 3/6~ : Yend pumpdown 3/6~
   * 3/7,8: leak test 

 * Xend
  * --(2/11-2/14: Xend NAB installation)--
  * --(2/21: Removing FC, Payload release)--
   * --(2/27: ETMX release)--
  * 2/23-3/10: Xend Pcal finalization 1
   * Periscope installation
   * Beam position by IR and Green
  * 3/5: ETMX aligned
  * 3/5-3/10: Xend Pcal finalization 2
  * 3/11,12 Xend WAB finalization
  * 3/19,20, Heatlink etc.
  * 3/22 Pumping down
   * Leak test
 
 * 3/11~ : Night work starts

 * 2/18-3/22 DRMI commissioning
 * 3/11-4/5 Y arm commissioning
 * 3/25-4/5 FPMI commissioning

==== Note ====

=== Weekend ===

==== Report of work ====
 * Type B: Drifts of relative yaw on SRM are still lasting, but getting slower.
 * Type Bp: GAS filter keystone was almost touching something. PR3 looks healthy.
  * PR2 has some broken circuit for sensors.
  * One of the magnet on PRM was off.
  * some works inside the chamber still remains.

 * Type A: Alignment of IR to ETMX, and to ETMY, and Green to both ETM were done on Friday.
 * Type A: Some ETMX LVDT signals showed a 584Hz oscillation. It will be traced today.

 * CAL: EX alignment was done.
  * EX cryopayload had some strange signal. Connection will be checked again.

 * PEM: Ground vibration measurement is being continued. Air Conditioner was not a cause of the vibration.



==== Work plan ====